OpenStudy (anonymous):

Solve 5n - 8 = 24. Answer 3.2 -3.2 6.4 -6.4

OpenStudy (anonymous):

Add 8 Then divide by 5

OpenStudy (anonymous):

32/5 = 6.4

OpenStudy (anonymous):

24+8 = 32/5